From GlueXWiki
Jump to: navigation, search

October 30, 2014 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. DAQ (Beni)
  2. Electronics (Fernando)
  3. CDC update (Mike)
  4. FDC update (Vlad, Lubomir)
  5. Commissioning Wiki Page (all)
  6. Other


Participants: Naomi (CMU) Eugene, Fernando, Simon, Dave, Tim, Chris, Nick, Mike, Sean, Beni, Cody, Sergey and Lubomir (JLab).

FDC chiller

- The chiller failed again on Tuesday (see plot attached): at 8:23am the Fluorinert temperature started increasing (blue), when it reached 25C the software interlock (green step function) stopped the LV. The temperature increased up to 48C then went down to 20C, by itself, by the time we got down in the Hall at 9:15am. The temperature change was real since it was followed by a change of up to 2 Pa in the gas pressure in the chambers (brown).

- Dave: the chiller showed initially "11: Reservoir High Temp" then " 21: Controller Error FLT T2 Sensor". After some troubleshooting we were not able figure out if the sensor or the controller has a problem, but another problem appeared: "04: CPRSR Overheat FLT". At this point, the facility people found that the compressor failed again. The original compressor failed and was replaced 2-3 months ago. Dave talked to the SMC representative: the Controller/Sensor error won't go by itself, most likely it will show up again after the compressor is fixed. Eugene agreed to order a new controller and a sensor. The parts will be here on Monday.

- Tim: facilities are replacing the compressor now. The representative from Hermitage Automation is coming at 2pm to look at the unit. We will continue troubleshooting once the compressor is replaced.

- Lubomir will discuss with Hovanes possibilities for powering off the chiller in case of a failure like this using epics.


- Cody made changes in the f125 firmware and now the DAQ is running without any crashes or errors!!! All the CDC and FDC crates are read out, but right now there's no information in the FDC crates.

- Beni: David Lawrence found that the BCAL and CDC information is shifted by one event. This was later fixed by including one additional reset in the readout list.

CDC update

- Mike: CDC HV channels that were put on 100nA and 1uA scale both tripped during the photon beam alignment. Now both HV channels were switched to 10uA scale.

- Mike has a GUI that monitors the CDC rates online, using independent trigger-less readout of the crates. Mike is ready to work on a similar GUI for the FDC.

FDC update

- The final results of the GlueX detector survey were given to Tim (attached) where FDC packages are shown individually. The biggest deviation of the package planes from vertical position is ~0.1 deg which is expected to have small contribution to the resolution.

- Vlad applied the reconstructed gain coefficients of the third package to the data from the EEL: these 3% (R.M.S.) corrections result in improvement of the strip resolution from 150 microns to 140 microns.


- Lubomir has a working plugin for the FDC online monitoring that works with the latest f125 data format.

- Simon (answering Eugene's question) about the CDC alignment: no improvement in the resolution so far, limited by the statistics and the fact that the cosmic trigger selects tracks at one side of the detector. Expect better results with BCAL trigger and more statistics.